# Generalists vs. Specialists: Evaluating Large Language Models for Urdu

### v2

@@ 5.2 Generation
We present the evaluation of the generalists and the specialists for the generation tasks in Table 3. We observe that the fine-tuned Llama-FT model achieves the highest scores in several tasks, notably in summarization and translation from Urdu to English. GPT-4 shows a consistent performance with its best results often appearing in the 6-shot setting. The mT5 model also demonstrates strong performance in tasks such as transliteration and paraphrasing, benefiting from its extensive multilingual training on translation tasks. We now discuss the performance on each generation task in detail.
@@
Question-Answering. For question-answering, GPT-4 achieves the highest score with a SQuAD-F1 of 72.78 in the 6-shot setting. Llama-3-8b-FT closely follows with a score of 70.42, indicating the effectiveness of its fine-tuning. mT5 also performs well with a score of 69.66. The consistent improvement of GPT-4 with increasing shots suggests its capability to use more context effectively.
@@
Summarization. In the summarization task, Llama-FT achieves the highest ROUGE-L score of 31.01, significantly outperforming Llama-3-8b and GPT-4. mT5 also shows strong performance with a score of 30.72. GPT-4’s best performance is in the 3-shot setting with a score of 23.35. The fine-tuning of Llama contributes to its superior performance in capturing and summarizing Urdu content effectively.
@@
Paraphrasing. For paraphrasing, mT5 achieves the highest SacreBLEU score of 11.98. This indicates the advantage mT5 has due to its massive multilingual pre-training. Llama-FT follows with a SacreBLEU score of 10.17. GPT-4’s best score is 4.01 in the 6-shot setting.
Figure 10: Comparison between mismatches in the words predicted by mT5 and GPT-4 for the transliteration task.
@@
Transliteration. In the task of transliteration, mT5 leads with a SacreBLEU score of 40.23, followed by Llama-FT at 37.95. GPT-4’s performance peaks at 32.38 with 3-shot learning. Figure 10 shows the words with the highest mismatches in the transliterated text. To count these mismatches, we first tokenize the transliterated sentences and then count the instances where the predicted word differs from the ground truth. Smaller words such as “ye,” “ke,” “aik,” and “wo” pose challenges for GPT-4, resulting in higher mismatch counts. In contrast, mT5 demonstrates lower mismatches for these words.
@@
Translation (en-ur). For English to Urdu translation, mT5 achieves the highest SacreBLEU score of 18.35, indicating its proficiency in translating English to Urdu. Llama-FT follows with a score of 14.63. GPT-4’s performance is consistent, with its highest score being 12.62 in the 3-shot setting. The superior performance of mT5 is likely due to the inclusion of high-quality Urdu data in the multilingual C4 corpus used for its pre-training Xue et al. (2021).
@@
Translation (ur-en). In Urdu to English translation, Llama-FT excels with a SacreBLEU score of 29.18, outperforming other models. Surprisingly, contrary to the results in en-ur translation, mT5 shows a lower SacreBLEU of 21.55 compared to Llama-FT. GPT-4’s highest score is 19.18 in the 6-shot setting.
@@
We compare human evaluation and LLM-based evaluation for summarization, paraphrasing, transliteration, English to Urdu translation, and Urdu to English translation. We select a subset of size 50 (as done by Khondaker et al., 2023a for Arabic) from the test dataset of each task. Since GPT-4 produces three outputs (0-shot, 3-shot, and 6-shot), we select the one with the highest SacreBLEU or ROUGE-L score. For human evaluation, two annotators (native Urdu speakers) are presented with anonymized outputs of Llama, Llama-FT, mT5 and the selected GPT-4. They are asked to rank them from one to four based on the criteria in Appendix G, allowing multiple models to have the same rank. We prompt GPT-4 and Llama with the same criteria, to assign a score to the outputs of the mentioned tasks, and then ranking is done based on this score. We compare the human ranking, GPT-4 ranking, and Llama ranking using Krippendorff’s alpha to determine the inter-rater reliability presented in Table 4.
Task A B C
Summarization 0.684 0.504 0.502
Paraphrasing 0.710 0.592 0.471
Transliteration 0.694 0.510 0.253
Translation (en-ur) 0.728 0.592 0.392
Translation (ur-en) 0.730 0.474 0.307
Table 4: Column A shows the Krippendorff’s alpha between annotator 1 and annotator 2, B shows the alpha between annotator 1, annotator 2 and GPT-4, and C shows the alpha between annotator 1, annotator 2 and Llama.
For each task, the Krippendorff’s alpha value for human evaluation exceeds 0.67, which, according to Krippendorff’s interpretation is sufficient for a tentative conclusion to be drawn. Table 4 illustrates that the inclusion of GPT-4’s evaluation or Llama’s evaluation significantly reduces the alpha values meaning that the annotations done by GTP-4 and Llama have a low degree of agreement with the human annotators. However there is a higher agreement between the GPT-4 rankings and human rankings as compared to the agreement between Llama rankings and human rankings. We also observe that according to human evaluation GPT-4 outperforms the other models for all the specified tasks. Figure 11 presents the number of times each rank was assigned to GPT-4 for Urdu to English translation. Appendix F contains presents the rank counts for the other tasks.
Figure 11: Rank counts for GPT-4 in human evaluation of Urdu to English translation.
@@ 6 Conclusion
In this paper, we present a comprehensive evaluation of generalist models and specialist models on 7 classification tasks and 6 generation tasks for Urdu NLP. Our evaluation covers prompting techniques such as few-shot, chain-of-thought reasoning as well as the fine-tuning of LLMs. We found that specialist models quantitatively outperformed generalist models on 12 out of the 13 tasks. The results highlight the importance of fine-tuning models to achieve higher performance in domain-specific applications in a low-resource setting. However, generalist models, such as GPT-4, showcased better performance in the human evaluation of the generation tasks, highlighting the importance of qualitative evaluation in accurately assessing model performance. We also performed a LLM based evaluation of the outputs of the models for the generation tasks. The low agreement between the rankings done by LLMs and human rankings shows that LLMS struggle when it comes to low-resource language understanding.
One avenue for future research is to explore other strong general-purpose models (e.g., GPT-4o, Claude Opus) and expand the scope of the evaluation to more Urdu NLP tasks. Additionally, using RAG to find examples from training dataset for few-shot prompting would be an interesting experiment to enhance the performance of generalist models. In conclusion, while specialist models currently hold an edge in task-specific performance, generalist models’ adaptability remains valuable, and continuous advancements in LLMs promise further improvements for low-resource language processing.
@@ 7 Limitations
While our study provides valuable insights into the performance of generalist and specialist models for Urdu NLP tasks, it is important to acknowledge several limitations. The question-answering and sentiment analysis datasets used for training the specialist models are translated from English to Urdu. This translation process can introduce inaccuracies that may affect the models’ performance. The evaluation is conducted on a subset of 1000 data points for each task. While this size is manageable and allows for a cost-efficient evaluation, it may not be fully representative of the model’s performance. Although we use multiple annotators and calculate inter-rater reliability using Krippendorff’s alpha, there is still a degree of subjectivity that may influence the results.
@@ Acknowledgments
We are grateful for the time and effort put in by our research intern, Muhammad Suhaib Rashid who annotated our data. We are also grateful to OpenAI for supporting our work through their Research Access Program.
@@ A.1 Models
XLM-Roberta-large is availabe on Hugging Face [^9] https://huggingface.co/FacebookAI/xlm-roberta-large under MIT license. mT5-large is available on Hugging Face [^10] https://huggingface.co/google/mt5-large under Apache-2.0 license. Llama-3-8b-Instruct is available on Hugging Face [^11] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct under llama3 license. GPT-4 is available under proprietary licence. All models used in this paper comply with their respective license.
@@ A.2 Datasets
Urdu IMDB sentiment analysis dataset OBDL license. Urdu NER dataset is availabe under MIT license. The abuse detection dataset by Akhter et al., 2020 and by Amjad et al., 2022, Sarcastic Tweets Dataset (Khan and Najeeb, 2023), fake news dataset by Amjad et al., 2020, Topic Classification dataset, and Universal Dependencies (Nivre et al., 2020) are available under CC BY 4.0.
UQuAD dataset is under CC0-1.0 while UQA (Arif et al., 2024) is under CC BY 4.0. XSUMUrdu (Munaf et al., 2023) summarization dataset is also under CC BY 4.0 license. Paraphrasing dataset is under MIT license. Dakshina dataset (Roark et al., 2020) for transliteration is under CC BY-SA 4.0 and OPUS-100 (Zhang et al., 2020) (Tiedemann, 2012) for translation is under GPL-3.0 license.
@@ Appendix B Model Size and Budget
We fine-tuned XLM-Roberta-large for classification tasks which has 550 million parameters. We fine-tuned mT5-large for generation tasks which has 1.2 billion parameters. Llama-3-8b-Instruct has 8 billion parameters and is fine-tuned for both generation and classification tasks.
@@ Appendix E Reproducibility and Hyperparameter
Appendix F Human Evaluation
For the summarization task, GPT was assigned a rank one by annotator one 41 times and by annotator two 34 times as shown in figure 12. In the paraphrasing task, figure 13 shows that GPT was chosen as the best model 42 times by annotator one and 39 times by annotator two. For transliteration, GPT was selected as the best model 40 times by annotator one and 38 times by annotator two as shown in figure 14. In the English to Urdu translation task, figure 15 shows that GPT was the top model 46 times for both annotators. For Urdu to English translation, GPT was rated as the best model 40 times by annotator one and 45 times by annotator two as presented in figure 11.
Figure 12: Rank counts for GPT-4 in human evaluation of summarization.
Figure 13: Rank counts for GPT-4 in human evaluation of paraphrasing.
Figure 14: Rank counts for GPT-4 in human evaluation of transliteration.
Figure 15: Rank counts for GPT-4 in human evaluation of English to Urdu translation.
For the summarization task, annotator one assigned Llama a rank of one 9 times, while annotator two assigned it 17 times. In the paraphrasing task, annotator one selected Llama as the best model 12 times, and annotator two chose it 15 times. For transliteration, Llama was selected as the best model 5 times by annotator one and 6 times by annotator two. In the English to Urdu translation task, annotator one ranked Llama as the top model 3 times, and annotator two ranked it 1 time. For the Urdu to English translation task, annotator one assigned Llama rank one 12 times, while annotator two assigned it 10 times.
For the summarization task, annotator one assigned Llama-FT a rank of one 7 times, while annotator two assigned it 12 times. In the paraphrasing task, annotator one selected Llama-FT as the best model 4 times, and annotator two chose it 3 times. For transliteration, Llama-FT was selected as the best model 13 times by annotator one and 21 times by annotator two. In the English to Urdu translation task, annotator one ranked Llama-FT as the top model 7 times and annotator two ranked it 10 times. For the Urdu to English translation task, annotator one favored Llama-FT 14 times, while annotator two favored it 15 times.
For the summarization task, mT5 was assigned a rank one by annotator one 4 times and by annotator two 11 times. In the paraphrasing task, mT5 was picked as the best model 4 times by both annotators. For transliteration, mT5 was selected as the best model 19 times by annotator one and 21 times by annotator two. In the English to Urdu translation task, mT5 was chosen as the top model 13 times by annotator one and 10 times by annotator two. For the Urdu to English translation task, mT5 was rated as the best model 18 times by annotator one and 15 times by annotator two. Figure 16 presents a detailed comparison of rank counts for GPT-4, Llama, Llama-FT and mT5 in human evaluation and LLM evaluation of various NLP tasks.
@@ Appendix G Prompts and Evaluation Criteria
Table LABEL:fig:classprompts contains the CoT prompts used for the classification tasks. The prompts for classification without CoT are the same, except that the reasoning is not generated, and only the label is required in the output. Table LABEL:fig:genprompts contains the prompts used for the generation tasks. Table LABEL:fig:eval-criteria presents the criteria given to human evaluators, GPT-4 and Llama for assessing the quality of outputs for summarization, paraphrasing, transliteration, English to Urdu translation, and Urdu to English translation. For each evaluation, the LLMs are asked to reason the answer to improve the scoring of the outputs.
Figure 16: Rank counts for GPT-4 in human evaluation of English to Urdu translation.
@@
Use chain of thought (cot) to reason your answer.
@@
Use chain of thought (cot) to reason your answer.
@@
Use chain of thought (cot) to reason your answer.
@@
Use chain of thought (cot) to reason your answer.
@@
Use chain of thought (cot) to reason your answer.
@@
Use chain of thought (cot) to reason your answer.
@@
Use chain of thought (cot) to reason your answer.

### v3

@@ 5.2 Generation
We present the evaluation of the generalists and the specialists for the generation tasks in Table 3 and Table 4. We observe that the fine-tuned Llama-FT model achieves the highest scores in summarization and translation from Urdu to English. On the other hand GPT shows a consistent performance across all the tasks with its best results often appearing in the 6-shot setting. The mT5 model also demonstrates strong performance in tasks such as transliteration and paraphrasing, benefiting from its extensive multilingual training on translation tasks. We now discuss the performance on each generation task in detail.
GPT’s Wins Llama’s Wins mT5’s Wins Llama-FT’s Wins
Task Annot. 1 Annot. 2 Annot. 1 Annot. 2 Annot. 1 Annot. 2 Annot. 1 Annot. 2
Summarization 41 34 9 17 7 12 4 11
Paraphrasing 42 39 12 15 4 3 4 4
Transliteration 40 38 5 6 13 21 19 21
Translation (en-ur) 46 46 3 1 7 10 13 10
Translation (ur-en) 40 45 12 10 14 15 18 15
AI Assistant 49 49 3 3 0 0 1 1
Table 4: We report the number of times each model is ranked 1 by each annotator for the generation tasks.
5.2.1 Quantitative Evaluation
@@
Question-Answering. All the models show similar performance on question-answering task. GPT achieves the highest score with a SQuAD-F1 of 72.78 in the 6-shot setting. Llama closely follows with a score of 71.65 in 3-shot setting. Llama-FT has a lower score of 70.42, indicating that the translated UQA dataset is not very effective in improving the performance of the model. mT5 also performs well with a score of 69.66. The consistent improvement of GPT with increasing shots suggests its capability to use more context effectively.
@@
Summarization. In the summarization task, Llama-FT achieves the highest ROUGE-L score of 31.01, outperforming Llama and GPT. mT5 also shows strong performance with a score of 30.72. Non-fine-tuned Llama outperforms GPT in all few-shot settings. GPT’s best performance is in the 3-shot setting with a score of 23.35 while Llama’s best is at 6-shot setting with a score of 26.17. The fine-tuning of Llama contributes to its superior performance in capturing and summarizing Urdu content effectively.
@@
Paraphrasing. For paraphrasing, mT5 achieves the highest SacreBLEU score of 11.98. This indicates the advantage mT5 has due to its massive multilingual pre-training. Llama-FT follows with a SacreBLEU score of 10.17. Llama’s best score is 6.40 at 3-shot setting outperforming GPT’s best score of 4.01.
@@
Transliteration. In the task of transliteration, mT5 leads with a SacreBLEU score of 40.23, followed by Llama-FT at 37.95. GPT’s performance peaks at 32.38 with 3-shot learning. Llama shows poor performace in this task with the highest score of 21.35 at 6-shot setting. Figure 10 shows the words with the highest mismatches in the transliterated text. To count these mismatches, we first tokenize the transliterated sentences and then count the instances where the predicted word differs from the ground truth. Smaller words such as “ye,” “ke,” “aik,” and “wo” pose challenges for GPT, resulting in higher mismatch counts. In contrast, mT5 demonstrates lower mismatches for these words.
Figure 10: Comparison between mismatches in the words predicted by mT5 and GPT for the transliteration task.
@@
Translation (en-ur). For English to Urdu translation, mT5 achieves the highest SacreBLEU score of 18.35, indicating its proficiency in translating English to Urdu. Llama-FT follows with a score of 14.63. GPT’s performance is consistent, with its highest score being 12.62 in the 3-shot setting. Llama under-performs with a least highest score of 6.56 with 3-shot prompting. The superior performance of mT5 is likely due to the inclusion of high-quality Urdu data in the multilingual C4 corpus used for its pre-training Xue et al. (2021).
@@
Translation (ur-en). In Urdu to English translation, Llama-FT excels with a SacreBLEU score of 29.18, outperforming other models. Surprisingly, contrary to the results in en-ur translation, mT5 shows a lower SacreBLEU of 21.55 compared to Llama-FT. GPT’s highest score is 19.18 in the 6-shot setting and Llama’s higest score is 13.11 in the 3-shot setting.
5.2.2 Qualitative Evaluation
AI Assistant. GPT outperforms all the models with 49 out of 50 wins as shown in Table 4. Llama-FT is ranked one 0 times according to both annotators and mT5 is ranked number one only 1 time. The low performance of fine-tuned models, such as Llama-FT and mT5, can be attributed to the fact that the fine-tuning dataset was translated from English.
Based on quantitative evaluation, GPT only performs the best in the question-answering task based on the quantitative evaluation provided in Table 3. However, Table 4 shows that GPT performs the best across all other tasks as well, according to qualitative evaluations by human annotators with more than 90% win-rate for most of the tasks. Figure 11 presents the number of times each rank was assigned to GPT for Urdu to English translation. Appendix F contains presents the rank counts for the other tasks.
Figure 11: Rank counts for GPT in human evaluation of Urdu to English translation.
Based on human evaluation LLama outperforms mT5 and even the fine-tuned Llama-FT on summarization, paraphrasing and AI assistant task. On transliteration and both translation tasks Llama-FT outperforms mT5. This discrepancy between quantitative and qualitative evaluations highlights an important issue: the translation of datasets from other languages may introduce inconsistencies or artifacts that impact the quantitative metrics, especially for tasks like summarization, paraphrasing, and story generation. This observation stresses the need for the creation of native Urdu datasets, rather than relying on translations of existing datasets. It also reflects the limitations and lower quality of the currently available Urdu datasets, which further exacerbates these challenges.
5.3 Factors Contributing to Improved Performance
Task-specific fine-tuning on Urdu datasets significantly improves the performance of LLMs on Urdu tasks, as evidenced by the superior results of fine-tuned Llama, XLM-R, and mT5 on both classification and generation tasks. Additionally, the inclusion of high-quality Urdu data in pre-training datasets, such as the clean multilingual C4 corpus, enhances downstream performance, with mT5 achieving higher SacreBLEU scores on transliteration, translation, and paraphrasing tasks compared to GPT and fine-tuned Llama. Providing more context in prompts also boosts performance, as shown by GPT’s improvement with increasing shots (0, 3, 6). Moreover, high-quality data for Indic languages like Hindi in pre-training improves performance on Urdu, especially for models like XLM-R, which excel in cross-lingual tasks, outperforming others in abuse detection, sarcasm detection, and fake news detection. The lack of native Urdu datasets poses a significant challenge for Urdu NLP, as most available datasets are translated from other languages.
@@ 5.3 Factors Contributing to Improved Performance
We compare human evaluation and LLM-based evaluation for summarization, paraphrasing, transliteration, English to Urdu translation, and Urdu to English translation, presented in Table 5.
Task A B C D
Summarization 0.684 0.504 0.502 0.624
Paraphrasing 0.710 0.592 0.471 0.565
Transliteration 0.694 0.510 0.253 0.603
Translation (en-ur) 0.728 0.592 0.392 0.515
Translation (ur-en) 0.730 0.474 0.307 0.510
AI Assistant 0.894 0.767 0.721 0.775
Table 5: A: Krippendorff’s alpha between annotator 1 and annotator 2. B: Alpha between both annotators and GPT. C: Alpha between both annotators and Llama. D: Alpha between both annotators and Claude.
For each task, the Krippendorff’s alpha value for human evaluation exceeds 0.67, which, according to Krippendorff’s interpretation is sufficient for a tentative conclusion to be drawn. Table 5 illustrates that the inclusion of GPT’s evaluation, Llama’s evaluation or Claude’s evaluation significantly reduces the alpha values meaning that the annotations done by GTP, Llama and Claude have a lower degree of agreement with the human annotators. For summarization, transliteration, Urdu-to-English translation, and AI assistant tasks, Claude has the highest agreement with humans, at 0.624, 0.604, 0.510, and 0.775, respectively. This showcases that the LLMs exhibit bias when the evaluated responses include their own generated content. For paraphrasing and Urdu-to-English translation tasks, GPT has the highest alpha values of 0.592 and 0.474, respectively. There is a higher agreement between the GPT rankings and human rankings as compared to the agreement between Llama rankings and human rankings.
@@ 6 Conclusion
In this paper, we present a comprehensive evaluation of generalist models and specialist models on 7 classification tasks and 7 generation tasks for Urdu NLP. Our evaluation covers prompting techniques such as few-shot, CoT reasoning as well as the fine-tuning of LLMs. We found that specialist models quantitatively outperformed generalist models on 12 out of the 14 tasks. The results highlight the importance of fine-tuning models to achieve higher performance in domain-specific applications in a low-resource setting. However, generalist models, such as GPT, showcased better performance in the human evaluation of the generation tasks, highlighting the importance of qualitative evaluation in accurately assessing model performance. We also performed a LLM based evaluation of the outputs of the models for the generation tasks. The low agreement between the rankings done by LLMs and human rankings shows that LLMs struggle when it comes to low-resource language understanding. Furthermore, the evaluation indicates a bias in LLMs when assessing responses that include their own outputs.
7 Future Work
One avenue for future research is to explore other strong general-purpose models (e.g., GPT-4o, GPT-4o1 and Claude) and expand the scope of the evaluation to more Urdu NLP tasks. Additionally, using Retrieval-Augmented Generation (RAG) to find examples from training dataset for few-shot prompting would be an interesting experiment to enhance the performance of generalist models. In conclusion, while specialist models currently hold an edge in classification tasks performance, generalist models’ adaptability and better performance in generation tasks remains valuable, and continuous advancements in LLMs promise further improvements for low-resource NLP.
@@ 8 Limitations
While our study provides valuable insights into the performance of generalist and specialist models for Urdu NLP tasks, it is important to acknowledge several limitations. The question-answering and sentiment analysis datasets used for training the specialist models are translated from English to Urdu. This translation process can introduce inaccuracies that may affect the models’ performance. Additionally, the lack of native, high-quality Urdu datasets poses a significant challenge. Translated datasets often fail to capture the linguistic and cultural nuances inherent to Urdu, which can impact both training and evaluation outcomes. The evaluation is conducted on a subset of 1000 data points for each task. While this size is manageable and allows for a cost-efficient evaluation, it may not be fully representative of the model’s overall performance. Although we use multiple annotators and calculate inter-rater reliability using Krippendorff’s alpha, there is still a degree of subjectivity that may influence the results.
@@ Acknowledgments
We are grateful for the time and effort put in by our research intern, Muhammad Suhaib Rashid who annotated our data and Mustafa Abbas who worked on creating Wiki-UQA dataset. We are also grateful to OpenAI for supporting our work through their Research Access Program.
@@ A.1 Models
XLM-Roberta-large is availabe on Hugging Face [^9] https://huggingface.co/FacebookAI/xlm-roberta-large under MIT license. mT5-large is available on Hugging Face [^10] https://huggingface.co/google/mt5-large under Apache-2.0 license. Llama-3-8b is available on Hugging Face [^11] https://huggingface.co/meta-llama/Meta-Llama-3-8B under llama3 license. GPT-4 is available under proprietary licence. All models used in this paper comply with their respective license.
@@ A.2 Datasets
Urdu IMDB sentiment analysis dataset OBDL license. Urdu NER dataset is availabe under MIT license. The abuse detection dataset by Akhter et al. (2020) and by Amjad et al. (2022), Sarcastic Tweets Dataset (Khan and Najeeb, 2023), fake news dataset by Amjad et al. (2020), Topic Classification dataset, and Universal Dependencies (Nivre et al., 2020) are available under CC BY 4.0.
UQuAD dataset is under CC0-1.0 while UQA (Arif et al., 2024a) is under CC BY 4.0. XSUMUrdu (Munaf et al., 2023) summarization dataset is also under CC BY 4.0 license. Paraphrasing dataset is under MIT license. Dakshina dataset (Roark et al., 2020) for transliteration is under CC BY-SA 4.0 and OPUS-100 (Zhang et al., 2020) (Tiedemann, 2012) for translation is under GPL-3.0 license. UrduAssistant (Khalil, 2023) dataset has a MIT license.
@@ Appendix B Model Size and Budget
We fine-tuned XLM-Roberta-large for classification tasks which has 550 million parameters. We fine-tuned mT5-large for generation tasks which has 1.2 billion parameters. Llama-3-8b has 8 billion parameters and is fine-tuned for both generation and classification tasks.
@@ Appendix E Reproducibility and Hyperparameter
AI Assistant - - 2 8 2 8
Appendix F Human and LLM Evaluation
Figure 12 presents the rank distribution for all models across all generation tasks, as evaluated by two human evaluators: GPT-4-Turbo, Llama-3-8b, and Claude 3.5 Sonnet.
@@ Appendix G Prompts and Evaluation Criteria
Table LABEL:fig:classprompts contains the CoT prompts used for the classification tasks. The prompts for classification without CoT are the same, except that the reasoning is not generated, and only the label is required in the output. Table LABEL:fig:genprompts contains the prompts used for the generation tasks. Table LABEL:fig:eval-criteria presents the criteria given to human evaluators, GPT and Llama for assessing the quality of outputs for summarization, paraphrasing, transliteration, English to Urdu translation, and Urdu to English translation. For each evaluation, the LLMs are asked to reason the answer to improve the scoring of the outputs.
Figure 12: Rank counts of models for all generation tasks.
@@
Use Chain-of-Thought (cot) to reason your answer.
@@
Use Chain-of-Thought (cot) to reason your answer.
@@
Use Chain-of-Thought (cot) to reason your answer.
@@
Use Chain-of-Thought (cot) to reason your answer.
@@
Use Chain-of-Thought (cot) to reason your answer.
@@
Use Chain-of-Thought (cot) to reason your answer.
@@
Use Chain-of-Thought (cot) to reason your answer.